OpenStudy (anonymous):

430 randomly selected light bulbs were tested in a laboratory, 224 lasted more than 500 hours. Find a point estimate of the proportion of all light bulbs that last more than 500 hours

OpenStudy (anonymous):

\[x = \frac{\sum x}{n} = \frac{224}{430} = 0.521 = 52.1\%\]